Please format answers to with EXCEL or to put in EXCEL. Thank you!

5. [EXCEL] Working capital: Laurel Electronics reported the following information at its annual meeting: The company had cash and marketable securities worth $1,235,455, accounts payables worth $4,159,357, inventory of $7,121,599, accounts receivables of $3,488,121, short-term notes payable worth $1,151,663, and other current assets of $121,455. What is the company's net working capital?

6. [EXCEL] Working capital: The financial information for Laurel Electronics referred to in Problem 5 is all at book value. Suppose marking to market reveals that the market value of the firm's inventory is 20 percent below its book value, its receivables are 25 percent below their book value, and the market value of its current liabilities is identical to the book value. What is the firm's net working capital using market values? What is the percentage change in net working capital?

7. [EXCEL] Income statement: The Oakland Mills Company has disclosed the following financial information in its annual reports for the period ending March 31, 2017: sales of $1.45 million, costs of goods sold of $812,500, depreciation expenses of $175,000, and interest expenses of $89,575. Assume that the firm has an average tax rate of 35 percent. What is the company's net income? Set up an income statement to answer the question.
Answer #1

5. Net Working Capital = Total Current assets - Total Current Liabilities

Current Assets
Cash and Marketable Securities $     12,35,455
Accounts Receivable $     34,88,121
Inventory $     71,21,599
Other current Assets $       1,21,455
Total Current Assets $ 1,19,66,630
Current Liabilities
Accounts Payable $     41,59,357
Short term notes payable $     11,51,663
Total Current Liabilities $    53,11,020
Net Working Capital $    66,55,610

6.

Current Assets
Cash and Marketable Securities $     12,35,455
Accounts Receivable $     26,16,091 =3488121*0.75
Inventory $     56,97,279 =7121599*0.8
Other current Assets $       1,21,455
Total Current Assets $    96,70,280
Current Liabilities
Accounts Payable $     41,59,357
Short term notes payable $     11,51,663
Total Current Liabilities $    53,11,020
Net Working Capital $    43,59,260

Change in Net Working Capital = ($4359260-6655610)/6655610 = -34.50%

7.

Income Statement
Sales Revenue $     14,50,000
Cost of Goods sold $       8,12,500
Gross Profit $       6,37,500
Depreciation Expense $       1,75,000
Operating Income $       4,62,500
Interest Expense $          89,575
Income before taxes $       3,72,925
Income tax expense @35% $       1,30,524
Net Income $       2,42,401